OpenStudy (juan1857):

Write the polar equation r=2csctheta in rectangular form.

OpenStudy (kittiwitti1):

...? I thought we did this one \(\text{already}\)?

OpenStudy (juan1857):

we did, oh yeah, forgot to write it down, lol

OpenStudy (kittiwitti1):

xD Well, if you want to save this problem and edit it with one you need help on, I can wait :P

OpenStudy (juan1857):

ok I appreciate it

OpenStudy (juan1857):

this is my new question: Write the polar equation r=2csctheta in rectangular form.

OpenStudy (kittiwitti1):

Okay, well here's a hint: \(\csc{\theta}=\frac{1}{?}\)

OpenStudy (juan1857):

um 1/sin?

OpenStudy (kittiwitti1):

Yup! And what is \(r\sin{\theta}\) again? :D

OpenStudy (juan1857):

y

OpenStudy (kittiwitti1):

Yes :D

OpenStudy (juan1857):

yeah i saw

OpenStudy (juan1857):

oh nvm, lol

OpenStudy (kittiwitti1):

Okay so this is how I did it. \(r=2\csc{\theta}\)So this gets us the following:\[r=2\times\frac{1}{\sin{\theta}}\rightarrow r=\frac{2}{\sin{\theta}}...\]So then you move the sine value over to the left, isolating the 2, and you get \(r\sin{\theta}=2\)

OpenStudy (kittiwitti1):

And you know what the formula \(r\sin{\theta}\) equals now, so... :)

OpenStudy (juan1857):

y=2 right?

OpenStudy (kittiwitti1):

\(\text{CONGRATULATIONS!!}\) You won the game :D

OpenStudy (juan1857):

AND @kittiwitti1 does it again!!!!. Thanks alot

OpenStudy (kittiwitti1):

No problem ^-^